OpenStudy (anonymous):

Solve the inequality negative four over five times p is greater than or equal to negative eightand choose the number line graph the represents the answer.

OpenStudy (anonymous):

Select one: A. Horizontal number line with nine markings, Center mark is marked ten with a bold dot. Line to the right of ten is bolded. Right and left ends of line are arrows pointing outward. B. Horizontal number line with nine markings, Center mark is marked negative ten with a bold dot. Line to the right of negative ten is bolded. Right and left ends of line are arrows pointing outward. C. Horizontal number line with nine markings, Center mark is marked negative ten with a bold dot. Line to the left of negative ten is bolded. Right and left ends of line are arrows pointing outward. D. Horizontal number line with nine markings, Center mark is marked ten with a bold dot. Line to the left of ten is bolded. Right and left ends of line are arrows pointing outward.

OpenStudy (anonymous):

I would really like it if someone would explain how they got their answer, so I could know how to get the answer for future refrence

OpenStudy (texaschic101):

-4/5p >= - 18 Is this the problem ?

OpenStudy (anonymous):

Yes, I need to know how to get the answer!

OpenStudy (anonymous):

p = 10

OpenStudy (texaschic101):

I wrote the problem wrong... -4/5p > = -8 divide both sides by -4/5 and remember to change the sign because you are dividing by a negative number p <= -8/(-4/5) p <= -8 * -5/4 p <= 40/4 reducing to 10 open circle on 10, shading to the left...so I believe that is D any questions ?

OpenStudy (texaschic101):

oops...closed circle on 10, shading to the left

OpenStudy (anonymous):

Awesome Thank you sooo very much! I understand it now, I didn't change the sign, so I got the wrong answer!

OpenStudy (texaschic101):

I , at first, wrote the problem wrong...that is why it took me so long...I wasn't getting any of your answer choices...I finally figured it out..lol.....glad I could help :)